Suppose you performed another titration on a different sample of old hydrogen peroxide. Give the data in the table below, what is the percentage of hydrogen peroxide in that solution? The molar mass of hydrogen peroxide is 34.0147 g/mol and the density of the solution is 1.000g/mol
volume of hydrogen peroxide solution (mL) 10.00mL
volume of 0.200 M KMnO4 dispensed in the tiration (mL) 14.72mL
a) 10.0%
b)2.50%
c)5.00%
d)7.25%
1
Expert's answer
2018-11-12T15:53:09-0500

C(H2O2)*V(H2O2) = C(KMnO4)*V(KMnO4)

C(H2O2) = 0.2mol/L*0.01472L/0/01L = 0.2944mol/L

m(H2O2) = n*M = 0.2944mol*34.0147g/mol = 10 g

a) 10%